OpenStudy (anonymous):

simplify the expression (-5p^)4

OpenStudy (lgbasallote):

-5p^? is there something missing here?

OpenStudy (anonymous):

sorry it was supposed to be in front of the 4

OpenStudy (lgbasallote):

so -5p^4?

OpenStudy (anonymous):

(-5p)^4

OpenStudy (lgbasallote):

ahh... here's a hint on how to solve this... \[\huge (ab)^c \implies a^c b^c\] for example... \[\huge (2x)^3 \implies 2^3 x^3 \implies 8x^3\] does that help?

OpenStudy (anonymous):

I dont see how we got 8x^3 from 2^3x^3

OpenStudy (lgbasallote):

2^3 = 2 * 2 2 = 8

OpenStudy (lgbasallote):

that was 2 * 2 * 2

OpenStudy (anonymous):

ok that makes sense

OpenStudy (lgbasallote):

good. so can you solve (-5p)^4 on your own?

OpenStudy (anonymous):

i came up with 625p^4

OpenStudy (lgbasallote):

right. good job

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome